Evaluate the integral. 17/12 tan+ 3t dt -7/12 O la
Answer #1

T/12 s Tan (st) at -T//2 Put 3t=u dt = 1/3 du Ę It t = 7/12 =) u = T/ It t = -172U=-174 TY Tlu → s Tan (u) du = ď sa tante)